[Brahms-l] Shift report 20040110 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 10 2004 - 08:27:56 EST
------------------------------------------------------------
Date: Sat Jan 10 08:27:56 2004, Shift: 00:00-08:00
Supervisor: sjs, with: 
Log book#15 pages 28 - 28

Shift Summary:
------------------------------------------------------------
Beam from 1:30 to 5:30.  However, we had multiple problems
during this time.  
Runs 9511,9512,9526,9527 are for pedestals.
Run 9513 (FS@4/A3450;MRS@90/B1050) trig6= 127K
On changing field settings after 9513 noticed D2 measured
current seems to be fixed at a value near 1500A.  Turned off,
then back on D2. Tried raising current above 1500A but
failed to get measured current reading to track to higher
current.  Since we still had decent beam, decided to continue
at a lower field setting and come back to the D2 problem 
between stored.
Run 9514 (FS@4/A1219;MRS@90/B1050) trig6=36K.  Run was
stopped when I noticed the event counts were not increasing
even though we still had beam.  Tracked problem to TPM1 partition following Problem FAQ of our Users Guide.  Initial
attempts to fix the problem failed, so decided to use remainder of store with a FS only run.  However, at this time we had a cooling problem in the pump room and all of the magnet supplies tripped.  After a little over an hour the pumps (and magnet supplies) were restored. Almost got a
new run started when the beam was dumped.

During the break between stored, had crew come out to fix D2. Feeling is that the supply was going to the set level,
but that the readout was faulty.  D2 can now be raised to
full current with the measured current tracking.  

Also went through the User Problem FAQ more carefull and
discovered that in addion to resetting the VME and receiver
boards,  one should also cycle the TPC power (switch on 
black panel in rack 7).  After doing this, recovered tpm1.

EVERYTHING IS NOW WORKING! ...on our end. RHIC is now showing red with an AGS transition power supply problem.

Other accomplishment:  I updated the General Shift Instructions with some 2004 specific stuff, including the current settings for pedestal runs and with a notice of the
very nice spelling checker that Hiro has added to e-log 
(hope you're reading this Hiro!).
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 10 08:28:38 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 10 2004 - 08:28:47 EST